# Harnessing the Power of Large Language Models for Empathetic Response Generation: Empirical Investigations and Improvements

### v1

@@ Abstract
Empathetic dialogue is an indispensable part of building harmonious social relationships and contributes to the development of a helpful AI. Previous approaches are mainly based on fine small-scale language models. With the advent of ChatGPT, the application effect of large language models (LLMs) in this field has attracted great attention. This work empirically investigates the performance of LLMs in generating empathetic responses and proposes three improvement methods of semantically similar in-context learning, two-stage interactive generation, and combination with the knowledge base. Extensive experiments show that LLMs can significantly benefit from our proposed methods and is able to achieve state-of-the-art performance in both automatic and human evaluations. Additionally, we explore the possibility of GPT-4 simulating human evaluators.
@@ 1 Introduction
Empathetic dialogue plays an essential role in building harmonious social relationships (Zech and Rimé, 2005). The task of empathetic response generation involves understanding the user’s experiences and feelings, and generating appropriate responses Keskin (2014); Rashkin et al. (2019). Using dialogue systems to provide empathetic responses has advantages such as easy access and no time constraints. Figure 1 shows an example of the empathetic dialogue from the benchmark dataset.
@@ 3.2 Preliminary Exploration
[colback=gray!15,colframe=gray!15,width=arc=1mm, auto outer arc, ] Task Definition + Guideline Instruction + Exemplars (optional) + Dialogue Context
@@ 3.3.2 Improvement via Two-stage Interactive Generation
[colback=gray!10,colframe=gray!10,width=arc=1mm, auto outer arc, left = 2mm, right = 2mm, top = 2mm, bottom = 2mm, enhanced,frame hidden, ] [The first stage]
“Don’t rush to reply yet, let’s think step by step. Based on the existing dialogue, what may be the user’s emotion, and according to his description, what may be the situation when he feels this way?”
[The second stage]
“Now combine your thoughts with the existing dialogue context and give your response.”
@@ 5.1 Preliminary Exploration Results
Table 1 shows the automatic evaluation results between LLMs and baselines. LLMs significantly outperform existing SOTA baselines and achieve a significant improvement on all automatic metrics, especially diversity. For Dist-1/2, LLMs achieve 51.8%[=(2.96-1.95)/1.95] and 92.7%[=(18.29-9.49)/9.49] improvement, which demonstrates a significant advantage of LLMs in diverse language expression (mainly unigrams and bigrams). In terms of BERTScore and BERT, LLMs achieve the average improvement of 2.1%[=(2.6+1.6+2.1)/3] and 26.95%[=(18.6+35.3)/2], respectively. This highlights the power of LLMs’ in-context learning capability that can be quickly applied to unseen specific tasks. In addition, we observe that the number of exemplars is positively correlated with diversity performance, which suggests the addition of examplars can influence the linguistic habits of LLMs.

### v4

@@ Abstract
Empathetic dialogue is an indispensable part of building harmonious social relationships and contributes to the development of a helpful AI. Previous approaches are mainly based on fine small-scale language models. With the advent of ChatGPT, the application effect of large language models (LLMs) in this field has attracted great attention. This work empirically investigates the performance of LLMs in generating empathetic responses and proposes three improvement methods of semantically similar in-context learning, two-stage interactive generation, and combination with the knowledge base. Extensive experiments show that LLMs can significantly benefit from our proposed methods and is able to achieve state-of-the-art performance in both automatic and human evaluations. Additionally, we explore the possibility of GPT-4 simulating human evaluators. Our code and data are available from https://github.com/27182812/LLM4ED.
@@ 1 Introduction
Empathetic dialogue plays an essential role in building harmonious social relationships (Zech and Rimé, 2005). The task of empathetic response generation involves understanding the user’s experiences and feelings, and generating appropriate responses Keskin (2014); Rashkin et al. (2019). Using dialogue systems to provide empathetic responses has advantages such as easy access and no time constraints Sharma et al. (2020). Figure 1 shows an example of the empathetic dialogue from the benchmark dataset.
@@ 5.1 Preliminary Exploration Results
Table 1 shows the automatic evaluation results between LLMs and baselines. LLMs significantly outperform existing SOTA baselines and achieve a significant improvement on all automatic metrics, especially diversity. For Dist-1/2, LLMs achieve 51.8%[=(2.96-1.95)/1.95] and 92.7%[=(18.29-9.49)/9.49] improvement, which demonstrates a significant advantage of LLMs in diverse language expression (mainly unigrams and bigrams). In terms of BERTScore and BLEU, LLMs achieve the average improvement of 2.1%[=(2.6+1.6+2.1)/3] and 26.95%[=(18.6+35.3)/2], respectively. This highlights the power of LLMs’ in-context learning capability that can be quickly applied to unseen specific tasks. In addition, we observe that the number of exemplars is positively correlated with diversity performance, which suggests the addition of examplars can influence the linguistic habits of LLMs.
@@ 6 Conclusion and Future Work
Acknowledgements
This work is supported by the National Key Research and Development Program (No. 2022YFF0902100) and National Natural Science Foundation of China (No. 62076081 and No. 61936010).